OpenStudy (anonymous):

Stuck on this question. Consider the function h(x) for which h(5)=-9 and h'(5)=7. Find f'(5) for the function f(x)=h(x)/x.

OpenStudy (anonymous):

\[f(x)=\frac{h(x)}{x}\] \[f'(x)=\frac{xh'(x)-h(x)}{x^2}\] plug in 5 and see what you get

OpenStudy (anonymous):

I got 3.2 but it says it is wrong did I plug in wrong?

OpenStudy (phoenixfire):

\[f^{\prime}(5)={5h^{\prime}(5)-h(5) \over 5^2}={{(5)(7)-(-9)} \over 25}={35+9 \over 25}={44 \over 25}=1.76\]

OpenStudy (anonymous):

ahhh I see what I did wrong ty!